Is perpendicular to 4y=3x-8 and passes through (-12,7)
Thank you!

Answer by mananth(16949)

-3 x + 4 y = -8
Find the slope of this line

4 y = 3 x -8
Divide by 4
y = 3/ 4 x + -2
Compare this equation with y=mx+b, m= slope & b= y intercept
slope m = 3/4

The slope of a line perpendicular to the above line will be the negative reciprocal -4/3
Because m1*m2 =-1
The slope of the required line will be -4/3

m= -4/3 ,point ( -12 , 7 )
Find b by plugging the values of m & the point in
y=mx+b
7 = 16 + b
b= -9
m= -4/3
The required equation is y = -4/ 3 x -9
m.ananth@hotmail.ca